OpenStudy (anonymous):

Find the vertex, value of p, axis of symmetry, focus, and directrix of the parabola. I will write the equation down shortly

OpenStudy (anonymous):

y-4=-1/2(x-8)^2

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@nathan917

OpenStudy (nathan917):

y - 4 = -1/2(x - 8)^2 y = -1/2(x - 8)² + 4 Vertex (8,4) <=x value opposite of inside parentheses and outside number for y a = -1/2 a = 1/(4p) -1. . . 1 ---- = ----- 2 . . .4p -4p = 2 p = -1/2 Focus (8,3.5) <=vertex plus p value onto y Directrix: y = 4.5 vertex y value - p which is - (-.5) or + .5 Axis of Symmetry: x = 8 <= x value of vertex

OpenStudy (nathan917):

Whats the point of tags if no one comes? But me :D

OpenStudy (nathan917):

What i mean is when you tag some one they should come.

OpenStudy (anonymous):

@nathan917 haha ok im here now

OpenStudy (nathan917):

:D.

OpenStudy (anonymous):

yay u gonna help

OpenStudy (anonymous):

@nathan917

OpenStudy (nathan917):

Whats your Question? id be glad to help.

OpenStudy (anonymous):

Simplify. Assume that all expressions are defined. (1/x)+(1/2x)/((x+4)/(x-2))

OpenStudy (anonymous):

@nathan917 im sorry my equation button thing on here isnt working